# Adaptive Contrastive Decoding in Retrieval-Augmented Generation for Handling Noisy Contexts

### v1

@@ Abstract
When using large language models (LLMs) in knowledge-intensive tasks, such as open-domain question answering, external context can bridge a gap between external knowledge and LLM’s parametric knowledge. Recent research has been developed to amplify contextual knowledge over the parametric knowledge of LLM with contrastive decoding approaches. While these approaches could yield truthful responses when relevant context is provided, they are prone to vulnerabilities when faced with noisy contexts. We extend the scope of previous studies to encompass noisy contexts and propose adaptive contrastive decoding (ACD) to leverage contextual influence effectively. ACD demonstrates improvements in open-domain question answering tasks compared to baselines, especially in robustness by remaining undistracted by noisy contexts in retrieval-augmented generation.
@@
Youna Kim1, Hyuhng Joon Kim1, Cheonbok Park2, Choonghyun Park1, Hyunsoo Cho3, Junyeob Kim1, Kang Min Yoo1 2 4, Sang-goo Lee1, Taeuk Kim5 1Seoul National University, 2NAVER Cloud, 3Ewha Womans University, 4NAVER AI LAB, 5Hanyang University {anna9812, heyjoonkim, pch330, juny116, sglee}@europa.snu.ac.kr {cbok.park, kangmin.yoo}@navercorp.com, chohyunsoo@ewha.ac.kr kimtaeuk@hanyang.ac.kr
@@ 1 Introduction
While large language models (LLMs) (Touvron et al., 2023; Achiam et al., 2023) achieve remarkable performance levels across diverse benchmarks, they sometimes struggle to generalize to knowledge-intensive tasks such as open-domain question-answering (QA; Chen et al., 2017) and may also fail to capture long-tail knowledge, leading to unfaithful output generation (Mallen et al., 2023; Kandpal et al., 2023). One common approach to address these issues is fine-tuning the model, but this results in a quadratic rise in computational demands as the size of the LLMs increases exponentially (Longpre et al., 2023). To address these issues, researchers have been investigating strategies to combine non-parametric knowledge with LLMs during response generation without explicit re-training. This approach leverages external information from knowledge bases and enhances the capability of the LLMs dynamically, ensuring that the information is both current and accurate.
@@
Early studies in this field attempt to append relevant context information to generate more accurate responses (Asai et al., 2023a). Especially, contrastive decoding (Li et al., 2023; Malkin et al., 2022; Liu et al., 2021) yields significant enhancement in various tasks by amplifying the influence of the given context at decoding step (Shi et al., 2023; Zhao et al., 2024). While such methods work well when context information is correct and faithful, in real-world scenarios, context information is not always correct and may contain some noisy and unfaithful information. For instance, if the retrieval system pulls in irrelevant or contradictory information, it could lead to incorrect responses (Wang et al., 2024; Wu et al., 2024; Yu et al., 2024). This highlights the necessity for a generation model that can gauge the appropriateness of the context by itself, being robust to noise and unfaithful data to ensure the output remains reliable (Yoran et al., 2024).
@@
To assess whether the existing contrastive decoding approaches can be utilized in practice, we extend the setting to where the gold-standard context is not guaranteed, specifically in the retrieval-augmented generation (RAG) framework (Yao et al., 2022; Shi et al., 2024; Izacard et al., 2023). In this paper, we demonstrate that existing context-aware contrastive decoding approaches experience performance drops in open-domain question answering, especially when the retrieved context is noisy. To address this issue, we propose adaptive contrastive decoding (ACD), adaptively weighting the contrastive contextual influence on the parametric knowledge, making it suitable for noisy context settings.
@@
Incorporating the distinction between contextual and parametric knowledge, our approach aims to mitigate the dominance of potentially noisy contextual information in model output. We control contrastive contextual influence based on context’s contribution to LLM’s uncertainty reduction, thereby minimizing its disruptive effect during decoding. Through in-depth experiments with three open-domain QA datasets, we demonstrate the potential of the proposed approach with increased overall performance. Moreover, ACD enhances the performance significantly on the noisy context scenario while minimizing performance degradation on the gold context scenario compared to the baselines.
@@ 2.1 Problem Formulation
At decoding time step $t$, given the input $x$ and preceding sequences $y_{<t}$, a pretrained auto-regressive LLM $\theta$ computes the logit $\mathbf{z}_{t}\in\mathbb{R}^{|V|}$, where $V$ is the vocabulary, for the $t$-th token. In the open-domain QA task, a question $q$ serves as the input $x$, and $\mathbf{z}_{t}$ relies solely on LM’s parametric knowledge. When both $q$ and the retrieved context $c$ are provided as $x$, the logit is denoted as $\mathbf{z}_{t}^{c}\in\mathbb{R}^{|V|}$.
@@ 2.2 Contrastive Decoding
In cases where context cannot be blindly trusted, directly following the context-augmented distribution can increase the risk of being misled. Thus, we adopt the approach of adding the contextual influence, which contrasts with the LLM’s parametric knowledge, to the parametric distribution $\mathbf{z}_{t}$. With the contrastive decoding objective, $\mathbf{z}_{t}^{c}$ and $\mathbf{z}_{t}$ are ensembled to reflect the influence of external context on LLM’s parametric knowledge at each decoding step $t$. The probability distribution $P_{\theta}(Y_{t}|x,y_{<t})$ is modified by weighted adjustment based on the difference between $\mathbf{z}_{t}^{c}$ and $\mathbf{z}_{t}$, as represented in the following equation.
@@ 2.3 Adaptive Weight on Contextual Influence
LLM’s uncertainty is expressed with the entropy $H(Y_{t})$ on its probability distribution $P_{\theta}(Y_{t}|x,y_{<t})$ (Huang et al., 2023; Kuhn et al., 2023). While $H(Y_{t})$ reflects how much uncertainty the model has based on its parametric knowledge under the given question, $H(Y_{t}^{c})$ is influenced by the external knowledge within the retrieved context $c$. Generally, when the context is added, the entropy decreases. However, if the context is noisy, irrelevant, or provides no information to answer the given question, it may contribute to increased uncertainty instead.
Intuitively, if the retrieved context provides informative cues for answering the question, then $H(Y_{t}^{c})$ is expected to be lowered compared to $H(Y_{t})$. Conversely, if the context is non-helpful or even confusing the model prediction, $H(Y_{t}^{c})$ in predicting the next token is likely to be higher. This scenario would be particularly evident when LLM knows the answer with low $H(Y_{t})$.
Considering the above scenarios, the motivation behind the adaptive weight $\alpha_{ACD}$ is to assign a relatively smaller weight in cases where the context increases uncertainty by being uninformative or confusing for LLM in answering the given question. Thus, the value of $\alpha_{ACD}$ is set as the proportion of uncertainty contributed by $H(Y_{t})$ relative to the total uncertainty when considering both $H(Y_{t})$ and $H(Y_{t}^{c})$:
@@
With $\alpha_{ACD}$, the full ACD is obtained, and the vocab $v$ with maximum probability is selected as the next token under the following distribution:
@@
Table 1: EM accuracy of full data (All) and subsets with gold ($\text{Subset}_{\text{Gold}}$) and noisy contexts ($\text{Subset}_{\text{Noisy}}$).
@@ 3.1 Experimental Settings
We conduct experiments on open-domain QA datasets, TriviaQA (Joshi et al., 2017), Natural Questions (NQ; Kwiatkowski et al., 2019), and PopQA (Mallen et al., 2022) with Wikipedia contexts [^1] Wikipedia dump from Dec. 2018..
We use auto-regressive language models, Llama2 (7B & 13B, Touvron et al., 2023), Llama3 8B [^2] https://github.com/meta-llama/llama3, and Mistral 7B (Jiang et al., 2023). Utilizing Contriever-msmarco (Izacard et al., 2022) as a retriever, the top-1 retrieved context is appended to each question.
@@
As fundamental baselines, regular greedy decoding has been employed in open-book ($\text{Reg}_{Opn}$) and closed-book ($\text{Reg}_{Cls}$) settings. We compare our method against existing context-aware contrastive decoding methods, including Context-Aware Decoding (CAD; Shi et al., 2023) and Multi-Input Contrastive Decoding (MICD; Zhao et al., 2024). MICD presents two methods, referred to as MICDF and MICDD, which offer fixed and dynamic $\alpha$, respectively. Similar to our approach, to leverage the burden of hyperparameter search and dependency on fixed $\alpha$, MICDD determines $\alpha$ in use by comparing the highest token probabilities with and without given context.
Other implementation details can be found in Appendix B.
@@ 3.2 Main Results
As shown in Table 1, ACD outperforms the baselines across all datasets and models, especially considering the full test data (All). Similar trends in performance are observed for Llama3 8B and Mistral 7B, as detailed in Appendix C. When analyzing the performance by dividing the data into two subsets based on whether the retrieved context is gold ($\text{Subset}_{\text{Gold}}$) or not ($\text{Subset}_{\text{Noisy}}$), ACD achieves either the best or second-best performance. $\text{MICD}_{D}$ demonstrates performance comparable to ACD on $\text{Subset}_{\text{Noisy}}$. However, it shows a significant drop on $\text{Subset}_{\text{Gold}}$, indicating a tendency to ignore gold context while handling noisy context. It is notable that both CAD and $\text{MICD}_{F}$ exhibit a significant drop in their performance under noisy conditions.
@@
We aim to analyze the model’s performance across various aspects, focusing specifically on its parametric knowledge. We estimate whether the model possesses relevant parametric knowledge for a given question based on its accuracy when only the question is provided. We consider two subsets under the following conditions: (1) Known-noisy: the model has parametric knowledge and noisy context is retrieved. (2) Unknown-gold: the model does not have parametric knowledge and gold context is retrieved.
From Figure 1, we observe that ACD outperforms the baselines in Known-noisy. Notably, two approaches with adaptively adjusted weight, ACD and $\text{MICD}_{D}$, perform well in Known-noisy, while other baselines show a relative strength in Unknown-gold. However, these baselines also experience significant performance drops in Known-noisy, indicating distraction by noisy context despite correctly answering when only the question is provided. In both cases, ACD demonstrates better performance compared to $\text{MICD}_{D}$, overall showing a tendency towards reliability. The results from other models are in Appendix D.
3.3 Analysis
Correlation between Adaptive Weight and Context Noisiness
@@ 3.3 Analysis
While other baselines rely on the fixed hyperparameter of weight $\alpha$, ACD and MICDD adjust $\alpha$ during the decoding step. It depends not only on the noisiness of the retrieved context but also on whether the model’s parametric knowledge contains an answer to the given question. To exclude cases that are not directly related to the analysis of how weight is adjusted based on context quality, we use the same subsets, Known-noisy and Unknown-gold.
@@
Figure 2: EM score vs. alpha values range from 0.0 to 1.0. Dashed line shows the performance with $\alpha_{ACD}$.
To assess whether adaptively modified $\alpha_{ACD}$ contributes performance improvement, we fix the value of $\alpha$ within a range $[0,1]$ and test on datasets used in the main experiment. In Figure 2, it can be observed that using a fixed $\alpha$ results in degraded performance compared to ACD. Increasing the value of alpha, which enhances the contextual influence on the output distribution, initially leads to a rise in the EM score. However, beyond a certain point, further increasing alpha results in a decline in the EM score. In scenarios with potential noisy context, a fixed alpha value may not ensure optimal performance. Therefore, employing an adaptive weight, $\alpha_{ACD}$, to adjust the impact of contextual knowledge based on entropy is crucial for improving overall performance.
As an additional analysis, we provide a case study of ACD to show how the value of entropy contributes to adjust $\alpha_{ACD}$ and further examine ACD on knowledge flipping task in Appendix F and Appendix G.

### v2

@@ Abstract
When using large language models (LLMs) in knowledge-intensive tasks, such as open-domain question answering, external context can bridge the gap between external knowledge and the LLMs’ parametric knowledge. Recent research has been developed to amplify contextual knowledge over the parametric knowledge of LLMs with contrastive decoding approaches. While these approaches could yield truthful responses when relevant context is provided, they are prone to vulnerabilities when faced with noisy contexts. We extend the scope of previous studies to encompass noisy contexts and propose adaptive contrastive decoding (ACD) to leverage contextual influence effectively. ACD demonstrates improvements in open-domain question answering tasks compared to baselines, especially in robustness by remaining undistracted by noisy contexts in retrieval-augmented generation.
**footnotetext: Corresponding author.
@@
Youna Kim1, Hyuhng Joon Kim1, Cheonbok Park2 3, Choonghyun Park1, Hyunsoo Cho4, Junyeob Kim1, Kang Min Yoo1 2 5, Sang-goo Lee1 6, Taeuk Kim7 * 1Seoul National University, 2NAVER Cloud, 3KAIST AI, 4Ewha Womans University, 5NAVER AI LAB, 6IntelliSys, Korea, 7Hanyang University {anna9812, heyjoonkim, pch330, juny116, sglee}@europa.snu.ac.kr {cbok.park, kangmin.yoo}@navercorp.com, chohyunsoo@ewha.ac.kr kimtaeuk@hanyang.ac.kr
@@ 1 Introduction
While large language models (LLMs) (Touvron et al., 2023; Achiam et al., 2023) achieve remarkable performance levels across diverse benchmarks, they sometimes struggle to generalize to knowledge-intensive tasks, such as open-domain question-answering (QA; Chen et al., 2017), and may also fail to capture long-tail knowledge, leading to unfaithful output generation (Mallen et al., 2023; Kandpal et al., 2023). One common approach to address these limitations is fine-tuning the model, but this results in a quadratic rise in computational demands as the size of the LLMs increases exponentially (Longpre et al., 2023). To overcome this, researchers have been investigating strategies to combine non-parametric knowledge with LLMs during response generation without explicit re-training (Asai et al., 2023a). This approach leverages external information from knowledge bases and enhances the capability of the LLMs dynamically, ensuring that the information is both current and accurate.
Figure 1: An illustration of adaptive contrastive decoding (ACD). Entropy ($H$) changes depending on context relevance, affecting the adaptive weight ($\alpha_{\text{ACD}}$). Noisy context leads the model to incorrectly answer "Diede De Groot" when employing regular greedy decoding. ACD applies context-based adjustments, enabling the correct answer, "Sloane Stephens," despite the noise.
@@
Early studies in this field attempt to append query-relevant context to generate more accurate responses. Especially, contrastive decoding (Li et al., 2023; Malkin et al., 2022; Liu et al., 2021) yields significant enhancement in various tasks by amplifying the influence of the given context at decoding step (Shi et al., 2023; Zhao et al., 2024). While such methods work well when context information is correct and faithful, in real-world scenarios, context information is not always correct and may contain some noisy and unfaithful information. For instance, if the retrieval system pulls in irrelevant or contradictory information, it could lead to incorrect responses (Wang et al., 2024; Wu et al., 2024; Yu et al., 2024). This highlights the necessity for a generation model that can gauge the appropriateness of the context by itself, being robust to noise and unfaithful data to ensure the output remains reliable (Yoran et al., 2024).
@@
To assess whether the existing contrastive decoding approaches can be utilized in practice, we extend the setting to situations where the gold-standard context is not guaranteed, specifically in the retrieval-augmented generation (RAG) framework (Yao et al., 2022; Shi et al., 2024; Izacard et al., 2023). In this paper, we demonstrate that existing context-aware contrastive decoding approaches experience performance drops in open-domain question answering, especially when the retrieved context is noisy. To address this issue, we propose adaptive contrastive decoding (ACD), adaptively weighting the contrastive contextual influence on the parametric knowledge, making it suitable for noisy context settings (Figure 1).
@@
Incorporating the distinction between contextual and parametric knowledge, our approach aims to mitigate the dominance of potentially noisy contextual information in model output. We control contrastive contextual influence based on context’s contribution to the LLM’s uncertainty reduction, thereby minimizing its disruptive effect during decoding. Through in-depth experiments with three open-domain QA datasets, we demonstrate the potential of the proposed approach with increased overall performance. Moreover, ACD enhances the performance significantly on the noisy context scenario while minimizing performance degradation on the gold context scenario compared to the baselines.
2 Related Works
Context-Augmented Generation
Approaches for context-augmented generation have been developed to enhance the model’s limited parametric knowledge by providing external knowledge, enabling more factual and contextually accurate responses during inference (Zhou et al., 2023; He et al., 2024). To sufficiently incorporate the information from the context in model generation, contrastive decoding approaches are applied to overwrite the model’s parametric knowledge with external knowledge (Shi et al., 2023; Zhao et al., 2024). These context-aware contrastive decoding methods to generate responses faithful to the given context show effective performance in summarization (See et al., 2017; Narayan et al., 2018), knowledge conflict (Longpre et al., 2022), and question answering with gold-standard contexts.
Robustness in RAG Frameworks
@@ 3.1 Problem Formulation
At decoding time step $t$, given the input $x$ and preceding sequences $y_{<t}$, a pretrained auto-regressive LLM $\theta$ computes the logit $\mathbf{z}_{t}\in\mathbb{R}^{|V|}$, where $V$ is the vocabulary, for the $t$-th token. In the open-domain QA task, a question $q$ serves as the input $x$, and $\mathbf{z}_{t}$ relies solely on the LLM’s parametric knowledge. When both $q$ and the retrieved context $c$ are provided as $x$, the logit is denoted as $\mathbf{z}_{t}^{c}\in\mathbb{R}^{|V|}$.
@@ 3.2 Contrastive Decoding
In cases where context cannot be blindly trusted, directly following the context-augmented distribution can increase the risk of being misled. Thus, we adopt the approach of adding the contextual influence, which contrasts with the LLM’s parametric knowledge, to the parametric distribution $\mathbf{z}_{t}$. With the contrastive decoding objective, $\mathbf{z}_{t}^{c}$ and $\mathbf{z}_{t}$ are ensembled to reflect the influence of external context on the LLM’s parametric knowledge at each decoding step $t$. The probability distribution $P_{\theta}(Y_{t}|x,y_{<t})$ is modified by weighted adjustment based on the difference between $\mathbf{z}_{t}^{c}$ and $\mathbf{z}_{t}$, as represented in the following equation.
@@ 3.3 Adaptive Weight on Contextual Influence
The LLM’s uncertainty is expressed with the entropy $H(Y_{t})$ of its probability distribution $P_{\theta}(Y_{t}|x,y_{<t})$ (Huang et al., 2023; Kuhn et al., 2023). While $H(Y_{t})$ reflects how much uncertainty the model has based on its parametric knowledge under the given question, $H(Y_{t}^{c})$ is influenced by the external knowledge within the retrieved context $c$. Generally, when the context is added, the entropy decreases (Kendall and Gal, 2017). However, if the context is noisy, irrelevant, or provides no information to answer the given question, it may contribute to increased uncertainty instead.
Intuitively, if the retrieved context provides informative cues for answering the question, then $H(Y_{t}^{c})$ is expected to be lowered compared to $H(Y_{t})$. Conversely, if the context is non-helpful or even confusing the model prediction, $H(Y_{t}^{c})$ in predicting the next token is likely to be higher. This scenario would be particularly evident when the model knows the answer with low $H(Y_{t})$.
Considering the above scenarios, the motivation behind the adaptive weight $\alpha_{ACD}$ is to assign a relatively smaller weight in cases where the context increases uncertainty by being uninformative or confusing for the model in answering the given question. Thus, the value of $\alpha_{ACD}$ is set as the proportion of uncertainty contributed by $H(Y_{t})$ relative to the total uncertainty when considering both $H(Y_{t})$ and $H(Y_{t}^{c})$:
@@
With $\alpha_{ACD}$, the vocab $v$ with maximum probability is selected as the next token under the following distribution:
@@
Llama3 8B $\text{Reg}_{Cls}$ 61.67 - - 28.34 - - 32.65 - -
$\text{Reg}_{Opn}$ 61.27 86.94 36.02 33.30 63.10 14.40 39.73 82.95 8.64
CAD 49.70 72.45 27.31 29.17 58.39 10.64 35.86 76.82 6.40
MICDF 61.01 85.40 37.00 27.62 51.89 12.22 37.99 77.12 9.85
MICDD 64.01 86.08 42.28 30.72 53.96 15.98 41.35 79.32 14.04
ACD 66.32 89.20 43.81 35.48 62.03 18.65 43.25 84.48 13.60
Mistral 8B $\text{Reg}_{Cls}$ 63.72 - - 29.64 - - 29.04 - -
$\text{Reg}_{Opn}$ 60.45 86.85 34.48 32.55 64.67 12.18 38.28 81.26 7.36
CAD 44.69 66.89 22.85 24.10 52.25 6.25 33.93 73.95 5.15
MICDF 63.33 88.43 38.62 31.80 61.10 13.22 36.58 76.00 8.23
MICDD 66.97 89.24 45.05 33.24 57.89 17.61 39.87 78.46 12.11
ACD 67.82 90.16 45.83 35.37 62.17 18.38 41.47 82.90 11.68
Table 1: EM accuracy of full data (All) and subsets with gold ($\text{Subset}_{\text{Gold}}$) and noisy contexts ($\text{Subset}_{\text{Noisy}}$). The highest score is in bold, and the second-best is underlined.
@@ 4.1 Experimental Settings
We conduct experiments on open-domain QA datasets, TriviaQA (Joshi et al., 2017), Natural Questions (NQ; Kwiatkowski et al., 2019), and PopQA (Mallen et al., 2022) with Wikipedia contexts. [^2] Wikipedia dump from Dec. 2018.
We use auto-regressive language models, Llama2 (7B & 13B, Touvron et al., 2023), Llama3 8B, [^3] https://github.com/meta-llama/llama3 and Mistral 7B (Jiang et al., 2023). Utilizing Contriever-msmarco (Izacard et al., 2022) as a retriever, the top-1 retrieved context is appended to each question.
@@
As fundamental baselines, regular greedy decoding has been employed in open-book ($\text{Reg}_{Opn}$) and closed-book ($\text{Reg}_{Cls}$) settings. We compare our method against existing context-aware contrastive decoding methods, including Context-Aware Decoding (CAD; Shi et al., 2023) and Multi-Input Contrastive Decoding (MICD; Zhao et al., 2024). MICD uses inputs with and without context, along with an additional input with adversarial context, to generate the output distribution. MICD presents two methods, referred to as MICDF and MICDD, which offer fixed and dynamic $\alpha$, respectively. Similar to our approach, to leverage the burden of hyperparameter search and dependency on fixed $\alpha$, MICDD also determines $\alpha$ dynamically. In $\text{MICD}_{D}$, $\alpha$ is assigned as the maximum token probability with context ($\text{max}P_{wc}$) if $\text{max}P_{wc}$ exceeds the maximum token probability without context ($\text{max}P_{woc}$); otherwise, it is calculated as $1-\text{max}P_{woc}$.
@@ 4.2 Main Results
As shown in Table 1, ACD outperforms the baselines across all datasets and models within the RAG framework, particularly when considering the full test data (All). When analyzing the performance by dividing the data into two subsets based on whether the retrieved context is gold ($\text{Subset}_{\text{Gold}}$) or not ($\text{Subset}_{\text{Noisy}}$), ACD achieves either the best or second-best performance. $\text{MICD}_{D}$ demonstrates performance comparable to ACD on $\text{Subset}_{\text{Noisy}}$. However, it shows a significant drop on $\text{Subset}_{\text{Gold}}$, indicating a tendency to ignore gold context while handling noisy context. It is notable that both CAD and $\text{MICD}_{F}$ exhibit a significant drop in their performance under noisy conditions.
@@
We aim to analyze the model’s performance across various aspects, focusing specifically on its parametric knowledge. We estimate whether the model possesses relevant parametric knowledge for a given question based on its accuracy in a closed-book setting ($\text{Reg}_{Cls}$). We consider two subsets under the following conditions: (1) Known-noisy: the model has parametric knowledge of the given question and noisy context is retrieved. (2) Unknown-gold: the model does not have parametric knowledge of the given question and gold context is retrieved.
From Figure 2, we observe that ACD outperforms the baselines in Known-noisy. Notably, two approaches with adaptively adjusted weight, ACD and $\text{MICD}_{D}$, perform well in Known-noisy, while other baselines show a relative strength in Unknown-gold. However, these baselines also experience significant performance drops in Known-noisy, indicating distraction by noisy context despite correctly answering when only the question is provided. In both cases, ACD demonstrates better performance compared to $\text{MICD}_{D}$, overall showing a tendency towards reliability.
@@
4.3 Analysis
Correlation between Adaptive Weight and Context Noisiness
While other baselines rely on the fixed hyperparameter of weight $\alpha$, ACD and MICDD adjust $\alpha$ during the decoding step. It depends not only on the noisiness of the retrieved context but also on whether the model’s parametric knowledge contains an answer to the given question. To exclude cases that are not directly related to the analysis of how weight is adjusted based on context quality and the model’s parametric knowledge, we use the same subsets, Known-noisy and Unknown-gold.
@@ 4.3 Analysis
Handling Knowledge Conflict
With a knowledge conflict QA dataset, NQ-swap (Longpre et al., 2022), we verify whether the two decoding methods with dynamic weight, ACD and $\text{MICD}_{D}$, can generate context-based responses without considering a conflicting context as a noisy context. The conflicting context in the NQ-swap dataset is constructed by replacing the answer entity span in the original gold context with a random entity of the same type. Figure 3 illustrates that ACD consistently exceeds the performance of $\text{MICD}_{D}$ across all models and achieves results comparable to open-book regular decoding. The results indicate that the ACD’s approach remains effective even in settings where the context is relevant to the question but contradicts the model’s parametric knowledge.
@@
Figure 4: EM across alpha values ranges from 0.0 to 1.0. The dashed line indicates EM score with $\alpha_{ACD}$.
To assess the impact of $\alpha_{ACD}$ on performance, we fix the value of $\alpha$ within a range $[0,1]$ and examine whether employing ACD is more effective than optimizing a fixed weight. In Figure 4, it can be observed that using a fixed $\alpha$ results in degraded performance compared to ACD. Increasing the alpha value, which enhances the contextual influence on the output distribution, initially leads to a rise in the EM score. However, beyond a certain point, further increasing $\alpha$ results in a decline in the EM score. In scenarios with potential noisy context, a fixed $\alpha$ value may not ensure optimal performance. Therefore, employing an adaptive weight, $\alpha_{ACD}$, to adjust the impact of contextual knowledge based on entropy is crucial for improving overall performance.